Is there any way someone can show me step by step how yo solve a problem? I just need the basic understanding of solving a problem.
Answer by wuwei96815(245)

There are basically two types of problems, word problems and math problems. This is a simplification of course, but at least it will get you started in the right direction.

The problem with word problems is that you have to interpret exactly what is being asked. Then, hopefully, convert it into an equation that can be solved.
For instance, if I asked you how far you can travel in 3 hours going 60 mph exactly what do I want to know? I want to know distance. The equation for distance is distance = speed x time. So, the distance that you are looking for is 60mph (speed) x 3 hours (time) or 180 miles.

In a straight math problem the words have been eliminated so you only have to do the math. Remember, you must be given at least one equation for every unknown. If you are trying to get the values of two unknown, such as x and y, you cannot get an answer unless you have two different equations that tie the two unknowns together in some relationship. If you have one unknown and one equation, the easiest way to solve the problem is to isolate the unknown and solve for it. For instance, solve for x if 2x-3=x+2.
2x-3=x+2
2x-x=2+3
x=5

I hope that this elementary explanation of solving problems helps you.